The 1.00 cm long second hand on a watch rotates smoothly. (a) What is its angular velocity? (b) What is the speed of the tip of the hand?
Answer #1

time period = 60 seconds

w=2*pi/60 = 0.10466 rad/s

speed=rw= 0.10466cm/s
Answer #3

time period ,T= 60 seconds

angular velocity,w=2*pi/T= 0.10466 rad/s

speed=r*w= 0.10466cm/s = 1.0466*10^-3 m/s